OpenStudy (anonymous):

If the principal P = $800, the interest I = $400, and time t = 5 years, find the following. Whats the rate? and whats the future value?

OpenStudy (anonymous):

u are familiar with the formula right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, hold on the question just hit me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the rate is I / rt right? it should look like this, 400/ (800) (12/5),

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

next you switch to this,

OpenStudy (anonymous):

Thanks ill try, my brain is dead by now i've being doing homework since this morning and wrting and reasearch...Oh boy I am ready to go to bed. ;0(

OpenStudy (anonymous):

you arent the only one.

OpenStudy (anonymous):

i'm trying to finish the spanish flvs course

OpenStudy (anonymous):

Great!, This is the first time I found this website, I figure to give it a try

OpenStudy (anonymous):

i'll ty to fugre it out, hold on

OpenStudy (anonymous):

I really apreciated

OpenStudy (anonymous):

Send me a private message if you can help me with that question i need another one.

OpenStudy (anonymous):

1.2% which = 12%

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

know the future value?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

sorry ok

OpenStudy (anonymous):

FV = 1,200

OpenStudy (anonymous):

its 5,500,000

OpenStudy (anonymous):

FV = P + I = 800 + 400

OpenStudy (anonymous):

im guessing i was a little off on my answer?

OpenStudy (anonymous):

srry, i g2g

OpenStudy (anonymous):

1,200 = 800 ( 1 + r ) ^5 -> 1 + r = fifth root (1.5)

OpenStudy (anonymous):

Thus r = 1 - fifth root (1.5) = 8.45%

OpenStudy (anonymous):

@trueseminole Don't worry, we're all learning here :)

OpenStudy (anonymous):

Thanks guys ;)

OpenStudy (anonymous):

@ChlorophyII, so the answer for that specific question will be 8.5% and 1200 (future value)?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok great! Thanks a lot

OpenStudy (anonymous):

@ChlorophyII sorry for taking advantage of your smart brain, but is because i am so tired i can barely think no more, i need 2 more questions would you help?

OpenStudy (anonymous):

This is one A car dealer will sell you a used car for $6,219 with $219 down and payments of $113.78 per month for 60 months. What is the APR? (Use the formula on page 156. Round each answer to the nearest tenth.)

OpenStudy (anonymous):

and this is another one....Ill pay you if i have too i just dont want to think no more .... Calculate the monthly finance charge for the following credit card transaction. Assume that it takes 10 days for a payment to be received and recorded and that the month is 30 days long. Assume 365 days in a year. (Round your answer to the nearest cent.) $3,000 balance, 21% rate, $1,500 payment, adjusted balance method

OpenStudy (anonymous):

Just open the new post, we finish it up in no time :)

OpenStudy (anonymous):

Where? i open it up?

OpenStudy (anonymous):

oh never mind, sorry is my first time here heeh

OpenStudy (anonymous):

By the rule, one question per post!

OpenStudy (anonymous):

Of course, sorry I didnt know the rules here. ;O)

OpenStudy (anonymous):

Also, include the formula with your post will help the solver resolve your questions much more efficiently :)

OpenStudy (anonymous):

I posted already